At the endof the 2000 baseball season, the New York Yankees and the Cincinnati Reds had won a total of 31 World Series. The Yankees had won 5.2 times as many World Series as the Reds. How many World Series did each team win?
Answer by edjones(8007)

5.2+1=6.2
31/6.2=5
The reds won 5 and the yankees won 5.2*5=26
.
Ed